# New Types of Hydrogenlike matter Composed of Electron(s) and Meson(s)

### v1

@@ Abstract
In the present work, we predict the existence of new types of hydrogenlike matter, including hydrogenlike atoms ($\pi^{+}e^{-}$, $K^{+}e^{-}$, $D^{+}e^{-}$), hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$) and hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$). By solving the Schrödinger equation, the binding energy of hydrogenlike atoms is obtained as $E_{n}=-\frac{1}{2n^{2}}$. For hydrogenlike molecular ions and molecules, the variational method is employed to calculate the binding energies, i.e., $E_{+}=-0.587$ and $E_{0}=-1.139$ for hydrogenlike molecular ions and molecules, respectively. And the bond lengths for hydrogenlike molecular ions and molecules are also calculated, whose values are $2.003$ and $1.414$, respectively. Here all the quantities are in atomic unit for convenience. In additon, the strong interaction between the two constituent mesons is considered in our calculations, where we find that its influence on the hydrogenlike molecular ions and molecules can be neglected. Hopefully, these types of matter would be observed in the future with the improvement of accuracy in the high energy physical experiments.
@@ I Introduction
The exploration of microscopic matter and their inner structure can increase our understanding of nature, which is always full of interests and challenges.
J. Dalton from England proposed the atomic theory in 1803, and opened a new era in modern chemistry. Dalton’s theory not only explained many chemical phenomena and chemical laws successfully, but also revealed their inner connections, transforming ancient naive atomic ideas into a scientific atomic theory. Soon after that, J. L. Gay-Lussac from France summarized experimental results of various gas reactions and drew an important conclusion: when gases react with each other, they often combine in simple ratios, and the change in volume of the combined gas also has a simple proportional relationship with the volumes of the reacting gases. Lussac’s conclusion leaded to a problem of half an atom, which contradicts the indivisibility of atoms in atomic theory and is also impossible in chemical reactions. In order to solve this problem, Amedeo Avogadro from Italy examined Lussac’s experiments and his academic dispute with Dalton, and proposed a kind of particle of a new divisible level-molecules. In nearly half a century, Avogadro’s molecular theory was not recognized by the academic community. As the rapidly development of modern inorganic and organic chemistry during this period, Italian scientist Stanislao Cannizzaro provided theoretical justification for the concept of molecules, organically integrating Dalton’s atomic theory with Avogadro’s molecular theory, and systematically proposing the scientific atomic-molecular theory, which was ultimately established in academic community.
At the turn of 19th and 20th centuries, the discovery of X-rays, radioactivity, along with the development of spectroscopy, the proposal of empirical formulas for spectral laws, and the electron-theoretical explanation of the Zeeman effect had changed the traditional concepts that atoms were indivisible and elements unchangeable. Scientists successively proposed several models of atomic structure, but they were all successively disproved by experiments. In 1911, Ernest Rutherford designed the $\alpha$ particles scattering experiment, and revealed that most $\alpha$ particles passed directly through the gold foil, while a small fraction were deflected at large angles of even almost $180^{\circ}$ Rutherford (1911). In order to explain this phenomenon, Rutherford proposed the nuclear structure model, whose core idea is that there is an extremely small atomic nucleus at the center of an atom, and all electrons revolve around the atomic nucleus. However, it was impossible to understand the stability of atoms, and to explain why excited atoms emit discrete spectral lines rather than continuous ones. So in 1913, Bohr proposed a new model of atomic structure, which is based on stationary state hypothesis, frequency rule, and quantization condition of angular momentum Bohr (1913a, b, c). In 1924, inspired by the wave-particle duality of light, Louis de Broglie proposed that physical particles such as electrons also exhibit wave-particle duality de Broglie (1925). This viewpoint directly leaded to the birth of quantum mechanics, which provides a theoretical foundation for understanding the structure of atoms.
About 20 years after the nuclear structure model was proposed by Rutherford, a new kind of neutral particle named neutron was observed by J. Chadwick Chadwick (1932), which was a milestone in the history of nuclear physics development. Then Heisenberg and Iwanenko proposed that nuclear is make of protons and neutrons Heisenberg (1932); Ivanenko (1932). After that, nuclear physics, as an independent research field, continued to develop in experiments, theories and applications.
However, another puzzle was how protons and neutrons could be tightly bound together in a nucleus with a size of $10^{-13}$ centimeters. In 1935, Hideki Yukawa proposed that there exists a strong binding force between protons and neutrons within the atomic nucleus by exchanging an unknown meson Yukawa (1935). This meson, called pion ($\pi$), has a mass of approximately $100-200$ MeV. It was not until 1947 that C. Powell et al. discovered the $\pi$ meson Lattes et al. (1947a, b). In the 1950s, with the development of high-energy accelerators, a large number of particles directly involved in strong interactions were discovered in accelerator experiments, and they had extremely short lifetimes. By the early 1960s, more than 100 “elementary” particles had been discovered in nature, which could be classified into two categories according to their interactions:
  • Hadrons: Particles directly participating in strong interactions, such as protons, neutrons, $\pi$ mesons, strange particles and a series of resonant particles.
  • Leptons: Particles do not directly involved in strong interactions but only in electromagnetic and weak interactions, such as electrons, muons and neutrinos.
High-energy physics experiments further revealed that hundreds of hadrons were not “fundamental” but had internal structures. In 1964, M. Gell-Mann and Zweig proposed the quark model Gell-Mann (1964); Zweig (1964a, b), regarding quarks as a new hierarchial level in the structure of matter. All the hadrons known at that time were composed of three fundamental quarks: the up quark (u), down quark (d), and strange quark (s). Numerous high-energy physics experiments in the 1960s confirmed the existence of quarks and the success of the quark model. Subsequent discoveries expanded the quark family. In 1974, Sounuel C. C. Ting and Richter discovered the charm quark (c) Aubert et al. (1974); Augustin et al. (1974). In 1977, Lederman et al. discovered the bottom quark (b) Innes et al. (1977). In 1995, the top quark (t) was discovered Abe et al. (1994a, b). The six quarks together with their corresponding antiquarks serve as the “building blocks” for all hundreds of hadrons. Meanwhile, the discovery of leptons also reached six types: electrons, electron neutrinos, muon, muon neutrinos, tau lepton, and tau neutrinos. Thus, quarks and leptons are currently recognized as the deepest and smallest constituent components of matter at our present level of understanding. With the further development of accelerator and detection technologies, nowadays many new particles have been observed including Higgs, the traditional mesons and baryons, tetraquark states ($Z_{c}(3900)$, $T_{cc}^{+}$, $\cdots$), and pentaquark states ($P_{c}(4312)$, $P_{c}(4380)$, $P_{c}(4440)$, $P_{c}(4457)$, $\cdots$), see Refs. Chen et al. (2016); Liu et al. (2019); Olsen (2015); Hosaka et al. (2016); Lebed et al. (2017); Esposito et al. (2017); Guo et al. (2018); Ali et al. (2017); Karliner et al. (2018); Albuquerque et al. (2019); Brambilla et al. (2020); Guo et al. (2020); Meng et al. (2023); Bicudo (2023); Liu et al. (2024); Chen et al. (2023); Dong et al. (2021, 2017); Yang et al. (2020); Drenska et al. (2010); Briere et al. (2016); Richard (2016); Mezzadri and Spataro (2022); Johnson et al. (2024) for reviews.
@@
Besides, some kinds of composite particles have also been studied, which involve few-lepton systems (see Ref. Ma et al. (2025) and the references there in), muonic atoms, hadronic atoms (see the reviews Friedman and Gal (2007); Batty et al. (1997); Gal et al. (2016)), and $\pi-\mu$ Coulomb bound states Coombes et al. (1976). In Refs. Abdel-Raouf et al. (1998); Rebane (2012a); Bressanini et al. (1998); Rebane (2012b); Richard (1994); Varga et al. (1997); Mitroy and Novikov (2004); Rebane (2012c); Bressanini et al. (1997); Rebane and Markovskiĭ (2001); Rebane and Markovski (2003); Gridnev and Greiner (2005); Van Hooydonk (2005); El-Khateeb (2007); Suzuki et al. (1998); Richard (2002); Armour et al. (2005); Emami-Razavi (2015); Emami-Razavi and Darewych (2021), four-body-exotic systems were discussed, including the predictions of $\pi^{+}\pi^{+}e^{-}e^{-}$ and $K^{+}K^{+}e^{-}e^{-}$ bound states Rebane (2012a, b). Hydrogenlike charm-pion or charm-kaon matter was predicted in Ref. Luo et al. (2023).
[…]
@@
This paper is organized as follows. After the introduction, the theoretical framework is described in Sec. II, which contains the calculations of $Me^{-}$, $MMe^{-}$ and $MMe^{-}e^{-}$ systems ($M$ = $\pi^{+}$, $K^{+}$, $D^{+}$). After that, we show the numerical results in Sec. III and give a discussion in Sec. IV whether the strong interaction affect the systems. In Sec. V, a short summary is given.
@@ II.2 Hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$)
Like the case of hydrogen molecular ion, we adopt the variational method to calculate the binding energy of the systems. In this work, we only focus on the ground eigenstate case. In light of that the electron is only affected by the attractive Coulomb force of the two identical mesons, it is naturally to choose the tried wave function as the superposition as follows
@@
Since the potential of the system is invariant when performing the reflection with respect to the midpoint of the two mesons, the state of the electron can be classified according to the reflection symmetry. Here we have $C_{a}=C_{b}\equiv C_{+}$ for even parity, and $C_{a}=-C_{b}\equiv C_{-}$ for odd parity. So the tried wave function can be expressed as
@@
±$\pm\big{[}\frac{\lambda^{3}R}{2}+\frac{\lambda^{2}}{2}-2\lambda^{2}R-2\lambda-\frac{\lambda^{4}R^{2}}{6}\big{]}e^{-\lambda R}\bigg{\}}+\frac{1}{R}.$
@@ II.3 Hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$)
In 1927, Heitler and London successfully applied quantum mechanics to study the structure of hydrogen molecule Heitler and London (1927), and pioneered the field of quantum chemistry.
@@
The distance $R$ between the two mesons is treated as a parameter in our calculation. And next we utilize the variational method to get the eigenenergy and wave function of the ground state. In view that when $R$ becomes very large, the ground-state wave function of the hydrogenlike molecule can approximately written as the product of the ground-state wave functions of the two hydrogenlike atoms, considering the exchange symmetry of the two electrons, the tried wave function can be chosen as
@@ III NUMERICAL RESULTS
Figure 2: (color online) The electron cloud of hydrogenlike atoms and molecular ions in $xy$ plane. (a), (b), (c) and (d) corresponds to $Me^{-}\ (1S)$, $Me^{-}\ (2S)$, $Me^{-}\ (3S)$ and $MMe^{-}e^{-}$ with $M=\pi^{+},K^{+},D^{+}$, respectively.
@@
Figure 3: The line shapes of the $E_{+}$, $E_{-}$, $E_{0}$ and $E_{1}$. The variational parameter $\lambda$ is fixed as 1.238 for $E_{+}$, $E_{-}$, and as 1.166 for $E_{0}$, $E_{1}$.
For hydrogenlike molecule, the spin wave function of the electron pair is antisymmetric, i.e., the two electrons are in a spin singlet, whereas the spatial wave function is symmetric. So the electrons can yet close to each other, which means they can form electron cloud with relatively higher density. And the electron cloud has a strong attractive force on the two mesons, such that the two hydrogenlike atoms combine together. The structure that the electron pair is shared by both mesons and their spins are in a singlet is so called covalent boud.
@@ IV Discussion: Does the strong interaction affect the systems?
The constant $g_{V}$ is determined from the experimental value of $g_{\rho\pi\pi}$, i.e., $g_{V}=5.80\pm 0.9$ Harada and Yamawaki (2003). Comparing $\mathcal{L}_{VPP}$ with the Lagrangian in Ref. Isola et al. (2003), we have
@@
[…]
Figure 4: The Feynman diagrams of the $\pi^{+}$-$\pi^{+}$, $K^{+}$-$K^{+}$ and $D^{+}$-$D^{+}$ interactions.
@@
of which the procedure of calculation is show in the appendix. Note that we have introduce the exponential form factor $F(q^{2})=e^{(q_{0}^{2}-\vec{q}^{2})/\Lambda^{2}}$ for each vetex in the diagrams. $\Lambda$ is the cutoff which is taken as 1956.95 (1000 $\rm{MeV}$ in the natural units) in the present work. The line shapes of the effective potentials in Eq. (51)-(53) are shown in FIG. 5.
@@ V Summary
In the present work, we propose three types of hydrogenlike structures by replacing the proton(s) in hydrogen atom, molecular ion and molecule with meson(s) of $\pi^{+}$, $K^{+}$ or $D^{+}$. In order to study these systems, we solve the Schrödinger equation for hydrogenlike atoms, and use the variational method for hydrogenlike molecular ion and molecule. Since in the hydrogenlike matter, the masses of the mesons are all much heavier than that of electron, just like the case of hydrogen atom, molecular ion and molecule, the structures and the binding energies are almost the same as those of hydrogen systems. For $\pi^{+}e^{-}$, $K^{+}e^{-}$ or $D^{+}e^{-}$ system, the binding energy is $E_{n}=-\frac{1}{2n^{2}}$, and the most probable radius is $r_{n}=n^{2}$. For the systems of $\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$ and $D^{+}D^{+}e^{-}$, there exist bound states if the spatial wave function is symmetric, while no bound states appear if the spatial wave function is antisymmetric. The binding energy of these hydrogenlike molecular ion is $-0.587$, and the bond length is $2.003$. For $\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$ and $D^{+}D^{+}e^{-}e^{-}$ systems, if the spin wave function of the electrons is antisymmetric, and the spatial wave function is symmetric, there exist bound states. In this case, the two electrons can get close to each other, and form electronic cloud with high density near the midpoint of the two mesons, which provides electromagnetic attraction on the two mesons. As a result, the systems combine together. The binding energy of these systems is $-1.193$, and the bond length is $1.414$. Note that if the spin wave function is symmetric while the spatial wave function is antisymmetric, there is no bound state.
The hydrogenlike matter predicted in the present work has not yet paid much attention, since it is challenging to search for it experimentally. However, we believe that these types of matter is still of much interest, and would be observed in the future with the improvement of accuracy in the high energy physical experiments.

### v2

@@ Abstract
In the present work, we predict the existence of new types of hydrogenlike matter, including hydrogenlike atoms ($\pi^{+}e^{-}$, $K^{+}e^{-}$, $D^{+}e^{-}$), hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$) and hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$). By solving the Schrödinger equation, the binding energy of hydrogenlike atoms is obtained as $E_{n}=-\frac{1}{2n^{2}}$. For hydrogenlike molecular ions and molecules, the variational method is employed to calculate the binding energies, i.e., $E_{+}=-0.587$ and $E_{0}=-1.139$ for hydrogenlike molecular ions and molecules, respectively. And the bond lengths for hydrogenlike molecular ions and molecules are also calculated, whose values are $2.003$ and $1.414$, respectively. Here all the quantities are in atomic units for convenience. In addition, the strong interaction between the two constituent mesons is considered in our calculations, where we find that its influence on the hydrogenlike molecular ions and molecules can be neglected. Comparisons of hydrogenlike molecular ion and molecule with the systems governed by the strong interaction are made, which suggests the possible existence of doubly heavy triquark, hidden heavy-flavor tetraquarks and doubly heavy tetraquarks. Hopefully, these types of matter would be observed in the future with the improvement of accuracy in the high energy physical experiments.
@@ I Introduction
The exploration of microscopic matter and its inner structure continues to be an area of great interest and challenge, deepening our understanding of nature. In particle physics, it is established that the material elements contain quarks, leptons, gauge bosons and Higgs boson, which together govern the microscopic world of particles and their production and decay properties. As the development of the high-energy physics experiments, a wide variety of hadronic states have been observed. These include conventional mesons and baryons, tetraquark states ($Z_{c}(3900)$, $T_{cc}^{+}$, $\cdots$), and pentaquark states ($P_{c}(4312)$, $P_{c}(4380)$, $P_{c}(4440)$, $P_{c}(4457)$, $\cdots$), see Refs. Chen:2016qju; Liu:2019zoy; Olsen:2014qna; Hosaka:2016pey; Lebed:2016hpi; Esposito:2016noz; Guo:2017jvc; Ali:2017jda; Karliner:2017qhf; Albuquerque:2018jkn; Brambilla:2019esw; Guo:2019twa; Meng:2022ozq; Bicudo:2022cqi; Liu:2024uxn; Chen:2022asf; Dong:2021bvy; Dong:2017gaw; Yang:2020atz; Drenska:2010kg; Briere:2016puj; Richard:2016eis; Mezzadri:2022loq; Johnson:2024omq for reviews.
@@
In addition, various types of composite particles have also been investigated, including few-lepton systems (see Ref. Ma:2025rvj and the references therein), muonic atoms, hadronic atoms (see the reviews Friedman:2007zza; Batty:1997zp; Gal:2016boi), and $\pi-\mu$ Coulomb bound states Coombes:1976hi; Aronson:1982bz. Refs. abdel1998existence; rebane2012existence; bressanini1998stability; rebane2012symmetry; richard1994stability; varga1997stability; mitroy2004stability; rebane2012stability; Bressanini:1997zz; rebane2001stability; rebane2003binding; gridnev2005proof; VanHooydonk:2004su; el2007investigation; Suzuki:1998wut; richard2002stability; armour2005stability; emami2015short; emami2021review discussed four-body-exotic systems, including the predictions of $\pi^{+}\pi^{+}e^{-}e^{-}$ and $K^{+}K^{+}e^{-}e^{-}$ bound states rebane2012existence; rebane2012symmetry. Hydrogenlike charm-pion or charm-kaon matter has also been predicted in Ref. Luo:2023hnp.
Inspired by these studies, we predict in this work new types of hydrogenlike matter composed of electron(s) and meson(s), i.e., $\pi^{+}e^{-}$, $K^{+}e^{-}$, $D^{+}e^{-}$, $\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $^{+}D^{+}e^{-}, D+D$$^{+}D^{+}e^{-}e^{-}$, as well as the previously studied $\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$ systems in Refs. rebane2012existence; rebane2012symmetry. For these composite systems, we not only investigate their bounding energies and the wave functions, but also discuss whether or not the strong interaction impact the bounding energies.
With ongoing progress in experimental facilities, the production of the hydrogenlike matter containing electrons and mesons may become possible. For instance, the proposed Super $\tau$-Charm Facility (STCF) Achasov:2023gey by the Chinese particle physics community may provide a potential platform for producing and studying these composite particles.
Besides the intrinsic value of exploring these electron-meson systems, the reason why these composite particles are of particular significance is that they are similar to the systems composed of heavy and light quarks, potentially offering insight into such configurations. For example, a tetraquark state contains two heavy quarks and two light antiquarks, while a hydrogenlike molecule comprises two electrons and two mesons, where the mesons are much heavier than the electrons. Thus, the hydrogenlike systems can be viewed as the QED analog of quark-based systems.
@@
This paper is organized as follows. After the introduction, the theoretical framework is described in Sec. II, which contains the calculations of $e^{-}, MM$$Me^{-}$ and $MMe^{-}e^{-}$ systems ($M$ = $\pi^{+}$, $K^{+}$, $D^{+}$). After that, we show the numerical results in Sec. III and give a discussion in Sec. IV whether the strong interaction affect the systems. In Sec. V, a short summary is given.
@@ II.2 Hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$)
Like the case of hydrogen molecular ion, we adopt the variational method to calculate the binding energy of the systems. In this work, we only focus on the ground eigenstate case. In light of that the electron is only affected by the attractive Coulomb force of the two identical mesons, it is naturally to choose the trial wave function as the superposition as follows
@@
Since the potential of the system is invariant when performing the reflection with respect to the midpoint of the two mesons, the state of the electron can be classified according to the reflection symmetry. Here we have $C_{a}=C_{b}\equiv C_{+}$ for even parity, and $C_{a}=-C_{b}\equiv C_{-}$ for odd parity. So the trial wave function can be expressed as
@@
$\pm\big[\frac{\lambda^{3}R}{2}+\frac{\lambda^{2}}{2}-2\lambda^{2}R-2\lambda-\frac{\lambda^{4}R^{2}}{6}\big]e^{-\lambda R}\bigg\}+\frac{1}{R}.$
@@ II.3 Hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$)
In 1927, Heitler and London successfully applied quantum mechanics to study the structure of hydrogen molecule Heitler1927, and pioneered the field of quantum chemistry.
@@
The distance $R$ between the two mesons is treated as a parameter in our calculation. And next we utilize the variational method to get the eigenenergy and wave function of the ground state. In view that when $R$ becomes very large, the ground-state wave function of the hydrogenlike molecule can approximately written as the product of the ground-state wave functions of the two hydrogenlike atoms, considering the exchange symmetry of the two electrons, the trial wave function can be chosen as
@@ III NUMERICAL RESULTS
Figure 2: (color online) The electron cloud of hydrogenlike atoms and molecular ions in xyxy plane. (a), (b), (c) and (d) corresponds to $Me^{-}\ (1S)$, $Me^{-}\ (2S)$, $Me^{-}\ (3S)$ and $MMe^{-}e^{-}$ with $M=\pi^{+},K^{+},D^{+}$, respectively.
@@
Figure 3: The line shapes of the $_{+}, EE_{-}$, $E_{0}$ and $E_{1}$. The variational parameter $\lambda$ is fixed as 1.238 for $_{+}, EE_{-}$, and as 1.166 for $E_{0}$, $E_{1}$.
For hydrogenlike molecule, the spin wave function of the electron pair is antisymmetric, i.e., the two electrons are in a spin singlet, whereas the spatial wave function is symmetric. So the electrons can yet close to each other, which means they can form electron cloud with relatively higher density. And the electron cloud has a strong attractive force on the two mesons, such that the two hydrogenlike atoms combine together. The structure that the electron pair is shared by both mesons and their spins are in a singlet is so called covalent bond.
@@ IV Discussion: Does the strong interaction affect the systems?
The constant $g_{V}$ is determined from the experimental value of $g_{\rho\pi\pi}$, i.e., $g_{V}=5.80\pm 0.9$ Harada:2003jx. Comparing $\mathcal{L}_{VPP}$ with the Lagrangian in Ref. Isola:2003fh, we have
@@
where the parameter $\beta$ is determined by the vector meson dominance Isola:2003fh; Colangelo:1993zq, i.e., $\beta=0.9$. $m_{P}$ is the mass of pseudoscalar charmed meson. The symbol $\langle\cdots\rangle$ is trace of a matrix.
Figure 4: The Feynman diagrams of the $\pi^{+}$-$\pi^{+}$, $^{+}-K+K^{+}$ and $^{+}-D+D^{+}$ interactions.
@@
of which the procedure of calculation is show in the appendix. Note that we have introduce the exponential form factor $F(q^{2})=e^{(q_{0}^{2}-\vec{q}^{2})/\Lambda^{2}}$ for each vertex in the diagrams. $\Lambda$ is the cutoff which is taken as 1956.95 (1000 $\rm{MeV}$ in the natural units) in the present work. The line shapes of the effective potentials in Eq. (51)-(53) are shown in FIG. 5.
@@ V Summary
In the present work, we propose three types of hydrogenlike structures by replacing the proton(s) in hydrogen atom, molecular ion and molecule with meson(s) of $\pi^{+}$, $K^{+}$ or $D^{+}$. In order to study these systems, we solve the Schrödinger equation for hydrogenlike atoms, and use the variational method for hydrogenlike molecular ion and molecule. Since in the hydrogenlike matter, the masses of the mesons are all much heavier than that of electron, just like the case of hydrogen atom, molecular ion and molecule, the structures and the binding energies are almost the same as those of hydrogen systems. For $\pi^{+}e^{-}$, $K^{+}e^{-}$ or $D^{+}e^{-}$ system, the binding energy is $E_{n}=-\frac{1}{2n^{2}}$, and the most probable radius is $r_{n}=n^{2}$. For the systems of $\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$ and $D^{+}D^{+}e^{-}$, there exist bound states if the spatial wave function is symmetric, while no bound states appear if the spatial wave function is antisymmetric. The binding energy of these hydrogenlike molecular ion is 0.587-0.587, and the bond length is $2.003$. For $\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$ and $D^{+}D^{+}e^{-}e^{-}$ systems, if the spin wave function of the electrons is antisymmetric, and the spatial wave function is symmetric, there exist bound states. In this case, the two electrons can get close to each other, and form electronic cloud with high density near the midpoint of the two mesons, which provides electromagnetic attraction on the two mesons. As a result, the systems combine together. The binding energy of these systems is 1.139-1.139, and the bond length is $1.414$. Note that if the spin wave function is symmetric while the spatial wave function is antisymmetric, there is no bound state.
Comparisons of hydrogenlike molecular ion and molecule with the systems governed by the strong interaction are made in the following. In the hydrogenlike molecular ion, the two mesons repel each other, while the electron is attracted to each meson. This configuration is analogous to the doubly heavy triquark system, where the two heavy quarks form a color sextet, and this diquark together with the light antiquark form a color triplet. In such a color structure, the two heavy quarks experience a repulsive interaction, whereas the light quark and each heavy quark are mutually attractive. The existence of hydrogenlike molecular ions may suggest that doubly heavy triquarks could also exist, even though they cannot be directly observed in high-energy experiments due to not being color singlets.
Similarly, for hidden-flavor tetraquarks, if both the heavy quark pair $Q\bar{Q}$ and the light quark pair are in a color octet, the forces between the two heavy quarks and between the two light quarks are repulsive, while the interactions between heavy and light quarks are attractive. This pattern of repulsions and attractions mirrors that of the hydrogenlike molecule. On the other hand, in doubly heavy tetraquarks where $QQ$ is in a color sextet and $\bar{q}\bar{q}$ is in an antisextet, the interaction pattern again resembles that of the hydrogen-like molecule. Therefore, our findings on hydrogen-like molecular systems may offer insights into the possible existence of both hidden heavy-flavor and doubly heavy tetraquarks.
The hydrogenlike matter predicted in the present work has not yet received much attention, since its experimental observation remains challenging. With the improvement of accuracy and the development of detection technology in high energy physical experiments, these types of matter may be observed in the future, where the possible processes ParticleDataGroup:2024cfk include $K_{L}^{0}\to\pi^{+}e^{-}\nu_{e}$ for $\pi^{+}e^{-}$ atom, $D^{0}\to K^{-}e^{+}\nu_{e}$ for $K^{-}e^{+}$ atom for instance.